OpenStudy (anonymous):

Find the value of 53³ using the identity (x+y)³=x³+3x²y+3x²y+y³ Hint:53³=(50+3)³; therefore, x=50 and y=3.

OpenStudy (bobo-i-bo):

What do you not understand?

OpenStudy (anonymous):

Everything

OpenStudy (anthonyym):

Well it says 53³ = (50+3)³ And we are using the identity (x+y)³=x³+3x²y+3x²y+y³. So (50+3)³ = (x+y)³, and x = 50, y = 3. Do you understand how x = 50 and y = 3?